OpenStudy (anonymous):

can we make a 3D box using latex? @Nnesha

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

http://openstudy.com/study#/updates/54ee31cae4b0dc552b64ba51 latex tutorialz

Nnesha (nnesha):

i guess but i dunno how

OpenStudy (anonymous):

i cannot see 3D box latex in http://openstudy.com/study#/updates/54ee31cae4b0dc552b64ba51

OpenStudy (anonymous):

Anyway,thnx @Nnesha

OpenStudy (anonymous):

@marc

OpenStudy (anonymous):

I mean @MARC_

OpenStudy (anonymous):

I asked how to make boxes in LaTeX...

OpenStudy (anonymous):

do u know how to make 3D box in latex? @AnswerMyQuestions

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

I don't think that's possible

OpenStudy (anonymous):

Hold on...

OpenStudy (anonymous):

I FOUND IT

OpenStudy (anonymous):

Made by @thomaster

OpenStudy (anonymous):

\(\newcommand{\starup}{\color{white}{^\star}} \newcommand{\stardown}{\color{white}{_\star}} \bbox[2.3pt,#002C77]{ \begin{array}{} \Large\starup\ \stardown\starup\ \stardown\starup\ \stardown\starup\ \stardown\starup\ \stardown\\ \Large\starup\ \stardown\starup\ \stardown\starup\ \stardown\starup\ \stardown\starup\ \stardown\\ \Large\starup\ \stardown\starup\ \stardown\starup\ \stardown\starup\ \stardown\starup\ \stardown\\ \Large\starup\ \stardown\starup\ \stardown\starup\ \stardown\starup\ \stardown\starup\ \stardown\\ \end{array}} \begin{array}{} \bbox[6pt,#CC0C2F]{\hspace{152pt}}\\ \bbox[5pt,#CC0C2F]{\hspace{154pt}}\\ \bbox[5pt,#CC0C2F]{\hspace{154pt}}\\ \bbox[5pt,#CC0C2F]{\hspace{154pt}}\\ \end{array}\\ \begin{array}{} \bbox[5pt,#CC0C2F]{\hspace{250pt}}\\ \bbox[5pt,#CC0C2F]{\hspace{250pt}}\\ \bbox[5pt,#CC0C2F]{\hspace{250pt}}\\ \end{array}\\\)

jaynator495 (jaynator495):

Yes Ive Litterally made one... -_-

jaynator495 (jaynator495):

Ive made a 3D box XD

OpenStudy (anonymous):

WOW. i can never do that never in my life... like.. legit.

OpenStudy (anonymous):

Thnx @AnswerMyQuestions

OpenStudy (anonymous):

:P